Find all the third root of( -8)?
(Complex numbers)
1
Expert's answer
2021-10-18T07:50:34-0400

Let X be the cube root of -8.

"X^3 = -8"

"X^3 + 8=0"

"(X+2)(X^2-2X+4)=0"

For X + 2 = 0; we have X = −2

"X^2-2X+4=0"

This is need to solve this using the quadratic formula:

"X=\frac{-(-2)\pm\sqrt{(-2)^2-4\times1\times4}}{2\times1}"

"X=\frac{2\pm\sqrt{4-16}}{2}"

"X=\frac{2\pm\sqrt{-12}}{2}"

"X=\frac{2\pm2\sqrt{-3}}{2}"

"X=1\pm\sqrt{-3}"

This means all the third root of( -8) are;

"-2, 1+i\sqrt{3}, 1-i\sqrt{3}"